OpenStudy (anonymous):

Find the sum and choose the correct answer. 2n^3 + 4n^2 - 7 and -n^3 + 8n - 9 -n^3 - 4n^2 - 8n - 16 n^3 + 4n^2 + 8n - 16 n^3 + 4n^2 + 8n + 16 3n^3 + 4n^2 + 8n-2

OpenStudy (anonymous):

4n+21+2n+21= -4 well first you've got to put teh like terms together so... 4n+2n+21+21= -4 now you have to solve it by ading the like terms so... 6n+42= -4 now you've got to do backtracking. yo always take the sign before a number with you while you do backtracking. so instead of doing 42 - 4 > you have to do -42 - 4 which gives you the answer of -46. so now it is 6n= -46 all you got to do now is 6 divided by -46 which gives you -7 and 4/6 therefore : n= -7 4/6 got it from yahoo answers